### v1

# Tactile Displays Driven by Projected Light

Tactile Displays Driven by Projected Light
Max Linnander1, Dustin Goetz1, Gregory Reardon1, Elliot Hawkes1, and Yon Visell1∗
1Department of Mechanical Engineering, Biological Engineering Program,
Media Arts & Technology Program,
University of California, Santa Barbara, USA.
Corresponding author. E-mail: yonvisell@ucsb.edu
Tactile displays that lend tangible form to digital content could profoundly transform how we interact with computers, much like visual displays have driven successive revolutions in computing over the past 60 years. However, creating tactile displays with the actuation speeds, dynamic ranges, and resolutions that are required for perceptual fidelity has proved challenging. Here, we present a tactile display that directly converts projected light into visible tactile patterns using an energetically passive, photomechanical surface populated with arrays of millimeter-scale optotactile pixels. The pixels transduce incident light into mechanical displacements through rapid, light-stimulated thermal gas expansion, yielding displacements of up to 1 millimeter and response times of 2 to 100 milliseconds. Our use of projected light for power transmission and addressing enables these displays to be scaled in size and resolution at sustainable cost and complexity. We demonstrate devices with up to 1,511 independently addressable pixels. Perceptual studies confirm the capacity of the display to accurately reproduce tactile patterns in location, timing, frequency, and structure. This research establishes a foundation for practical, versatile high-resolution tactile displays driven by light.
Introduction
Tactile displays give physical form to digital content, conveying information through tactile patterns that can represent data, user interfaces, or virtual objects. They could profoundly transform human-computer interaction, much like visual displays have driven successive revolutions in computing, from the first widespread computers of the 1970s to wearable and virtual reality systems of today  (?). Applications of tactile displays span many fields, including automotive interfaces that emulate physical controls with software reconfigurability  (?, ?, ?), educational tools such as animated haptic books  (?, ?), and user interfaces for virtual and augmented reality  (?, ?). However, creating multipurpose tactile displays with perceptual fidelity imposes stringent requirements on spatial resolution ($\lesssim$ 3 mm), response time ($<$50 ms), displacement ($>$ 0.2 mm), and force ($>$ 5 mN), with proportional output  (?, ?, ?, ?). Achieving these requirements in displays of useful size at sustainable costs will require significant advancements in actuation technologies, scalability, integration density, and fabrication methods.
A variety of materials and actuation methods have been studied for tactile display engineering, including electrostatic, piezoelectric, electromagnetic, fluidic, and electroosmotic methods  (?, ?, ?, ?, ?, ?, ?). However, the state of the art remains dominated by bespoke displays with a hundred or fewer pixels; see Table S1 and survey  (?). [^1] Braille graphics devices could be cited as exceptions, but are generally designed to represent static graphics or text. They use binary actuation with long refresh times, thus do not meet performance requirements for general purpose tactile display  (?). A major challenge is scaling electronic tactile displays to larger sizes and resolutions at sustainable cost and complexity while maintaining performance parameters. As size and resolution increase, the complexity of driving and control circuits grow rapidly. Moreover, modifying the configuration of such systems often requires electronic redesign, hindering development.
An intriguing alternative is to exploit light transmission for tactile actuation. Tactile displays are often designed to integrate with video displays that utilize optical transmission. Harnessing light for tactile actuation presents a synergistic opportunity that could offer practical and performance benefits. Optical energy can be modulated at high frequencies and directed with high spatial precision using mature technologies, including video and laser projection, enabling rapid and localized wireless energy delivery. Even compact laser modules can deliver substantial power via light (e.g., 5 W), as much as would be required to drive dozens of miniature vibration actuators  (?, ?). Utilizing projected light for tactile displays may simplify power distribution and control, reducing system complexity, and supporting scalability.
Despite this potential, the use of targeted light transmission for tactile actuation has been sparsely investigated. Prior studies have employed thermal mechanisms that convert absorbed optical energy into forces and displacements. Examples include solid thermoelastic expansion  (?), phase transitions in liquid crystal elastomers  (?, ?, ?), and liquid vaporization  (?). In phase change mechanisms, a substantial portion of input energy is consumed to overcome latent heat barriers, delaying actuation. Furthermore, thermal processes often yield slow response times – on the order of seconds or more – due to constraints imposed by heat transfer (Table S2). The characteristic timescale for heat transfer through an actuator of dimension $L$ is of order $\tau=L^{2}/\alpha$, where $\alpha$ is thermal diffusivity. For polymers and liquids, $\alpha$ is on the order of 10-7 m2/s. In contrast, gases possess thermal diffusivities that are two orders of magnitude higher, because diffusivity is inversely proportional to density. For this reason, gases may enable sufficiently fast refresh times to enable dynamic tactile displays driven by projected light.
Building upon this insight, we present a tactile display that uses projected light to activate tactile patterns across a surface populated with small scale ($L$ = 3 mm) gas-filled pixels. Pixels are activated through the photostimulated heating and transient expansion of the encapsulated gas (Fig. 1). Light pulses with energies of 2.5 to 125 mJ are sufficient to activate localized displacements of up to 1 mm and forces of up to 55 mN, with response times of 2 to 100 ms. The actuation method, optical addressing technique, and simplicity of fabrication make this technique practical and scalable. We present implementations with up to 1,511 independently addressable pixels distributed over a $15\times 15$ cm2 area – several times more pixels than any previous tactile display approximating the requirements cited above (Fig. 1H, Table S1). We show that these light-driven tactile displays are capable of representing a wide variety of spatiotemporal tactile patterns that can be accurately perceived.
Figure 1: Fig. 1. Tactile display driven by projected light. A) Projected light excites tactile patterns, consisting of mechanical signals produced by an array of optotactile pixels. B) Light is converted to forces and displacements via an assembly of simple patterned layers, easing manufacturing and scalability. From top: Elastic membrane (EcoFlex 00-10), cavity walls (polysiloxane, PS), thin photoabsorber (PGS, thickness: 17 $\mu$m), cavity walls (PS), optically transparent layer (acrylic). C) Cross-section view of pixel operation. An incident light pulse is converted into heat by the photoabsorber. Heat is transferred to air in the cavity, increasing gas temperature $T_{gas}$ and pressure $P$. Gas expansion drives the deflection of the elastic membrane. D) Finite element analysis of pixel thermal and mechanical response. E) Flexible tactile display with 357 pixels. Scale bar: 1 cm. F) Photostimulation yields both tactile and visual feedback, due to the transmission of scattered light through the membrane. Extended exposure image of 1,511 pixel display. Scale bar: 1 cm. G) The same technique can be used to realize displays that vary in dimensions and pixel count. From upper left: 1, 8, 357, and 1511 pixels. Scale bar: 1 cm. H) Comparison with prior tactile displays. The projected light tactile display in this work yields fast response rates, scales to pixel counts exceeding prior displays, and is uniquely capable of both visual and tactile display. Driving mechanisms: L = light, EM = electromagnetic, PN = pneumatic, ES = electrostatic, T = thermal, PZ = piezoelectric, C = combustion, and EO = electroosmosis. (See also Fig. S16, Table S1).
Design and Photomechanical Transduction Principle
The display functions as an interface that translates projected light into tactile mechanical signals, via an array of optotactile pixels (Fig. 1A). Each pixel operates on a thermodynamic principle, transducing heat from an absorbed light pulse into mechanical energy. Light entering the pixel via an optical window is converted to heat by a thin photoabsorber, which rapidly increases in temperature (Fig. 1C). Heat is transferred to gas within the cavity, causing a pressure increase dictated by the ideal gas law $PV=nRT_{gas}$. Gas expansion drives the deflection of an elastic membrane sealing the pixels, yielding localized mechanical forces $F(t)$ and displacements $d(t)$ that can be felt by touching the membrane (Fig. 1D).
This general strategy may be used to realize a large variety of displays. The following design choices exemplify one way it can be implemented. A multi-layered architecture was used to facilitate ease of manufacturing and scalability (Fig. 1B). Transparent acrylic forms a base layer, functioning as the optical window (Fig. S2). The base is adhered to two polysiloxane (PS) layers with arrays of cylindrical cavities (3 mm diameter, 1 mm height) (Fig. 1C, S4-5). Between these layers is a thin pyrolytic graphite sheet (PGS, thickness: 17 $\mu$m) that is patterned to form an array of photoabsorbers, which are suspended within each cavity by narrow bridge features (Fig. 1C-D,S4). PGS was selected due to its high thermal conductivity, optical absorptance ($\varepsilon$=65%; Fig. S2), and thermal stability [^2] PGS is stable up to 3000 C; Sustained temperatures over 400 C may favor the use of a more inert gas, such as nitrogen or helium, to inhibit oxidization.. Capping this assembly is the tactile surface – an elastomer membrane that seals the air-filled cavities. The display can be made flexible through the use of a thinner, or more elastic, base layer (Fig. 1E). The entire layered assembly is energetically passive.
The surface is stimulated by brief (1 to 50 ms) low-energy light pulses (2.5 to 125 mJ) delivered from a compact diode laser ($\lambda=450$ nm; Fig. 1C,S6). Scanning is facilitated by commodity x-y galvanometer mirrors, yielding spatiotemporal tactile patterns that can be felt, and also seen, due to the transmission of scattered light through the membrane (Fig. 1F). The mechanism of actuation, optical addressing technique, simplicity of fabrication, and use of widely available materials render this technology practical and scalable. To demonstrate the ease of manufacturing and scalability, we fabricated devices with 1, 8, 357, and 1,511 addressable pixels (Fig. 1G). The pixel dimensions (diameter 3 mm, height 1 mm) were chosen based on theoretical analyses and finite-element numerical studies (Text S1, S2). As demonstrated below, our technique achieves forces, displacements, resolutions, and response rates meeting perceptual requirements for tactile display, at display scales suitable for diverse applications (Fig. 1H, Fig. S16, and Table S1).
Thermo-Mechanical Response Under Photostimulation
Figure 2: Fig. 2. Thermal-mechanical response characterization. A,B) Absorber surface temperature and membrane deflection increased monotonically during illumination with power $P_{L}=2.5$ W over 50 ms. (Scale bar = 0.5 mm). C) Reducing the thermal bridge width from $w=$ 0.75 mm to 0.2 mm yielded higher temperatures and larger displacements. Scale bar: 0.5 mm. D) Experimental absorber temperature $T(t)$ and membrane displacement $d(t)$ vs. time $t$ for $w=$ 0.2, 0.25, 0.4, 0.55, and 0.75 mm. As $w$ decreased, more heat was retained in the absorber, yielding higher temperatures and larger displacements. For $w=0.2$ mm, at $t=$ 50 ms we observed $T=527^{\circ}$ C and $d=0.97$ mm. Grey dashed line: Numerical results for $w=0.25$ mm; see Fig. S1. E) The thermal resistance $R$ of the bridge increased as width $w$ decreased. The data agreed with the theoretically predicted relationship $R=a/w,$ with $a=$ 73.3 mm K/W, $r^{2}=0.98$. F) Peak membrane displacement $d$ and absorber thermal relaxation time $\tau$ increased with thermal bridge resistance $R$, while the relaxation rate $1/\tau$ decreased, indicating that tuning $w$ involves a tradeoff between displacement and relaxation rate.
We experimentally characterized the thermal and mechanical response of individual pixels under pulsed photostimulation at constant power $P_{L}$ for pulse length $t_{p}=50$ ms. We observed the mean temperature $T(t)$ in the photoabsorber to rise monotonically during exposure (Fig. 2A-D), driving expansion of air within the pixel cavity and deflection of the elastic membrane (Movie S1). Temperature and displacement both reached their maximum values near the time at which the optical pulse ceased, indicating a nearly static relationship between the absorber and air temperatures during heating (Fig. 2D, Fig. S8). Following the pulse, temperature and displacement relaxed toward baseline conditions.
We studied the effect of photoabsorber geometry by varying the width, $w$, of the bridges suspending it in the cavity, from $w=$ 0.2 to 0.75 mm (Fig. 2A-D). Larger absorber temperatures, $T(t),$ and displacements, $d(t),$ are obtained from pixels with smaller widths, $w$, indicating that bridge width has a dominant effect on heat transfer. We confirmed this using theoretical and numerical analyses of heat transfer (Text S1-2). The highest absorber temperatures and displacements were obtained at the smallest width, $w=0.2$ mm. Temperature increased by $507$ C and displacement reached a maximum of $0.97$ mm after 50 ms of photostimulation at 2.5 W. The thermal and mechanical response of the pixels can thus be tuned by varying the bridge width, $w$.
We analyzed the dynamic thermal response using the experimental data in tandem with a standard description of heat transfer (Text S1). An effective description of heat transfer indicates that the time scale $\tau$ governing the absorber temperature $T(t)$ response to photostimulation is given by $\tau=RC,$ where $C$ is the absorber heat capacity and $R$ is the effective thermal resistance of the absorber. We determined $R$ for each width $w$ by fitting the experimental data at the onset of photostimulation (Table S3). The effective thermal resistance was inversely proportional to bridge width, in agreement with the theoretically expected relationship $R=L(k_{xy}hw)^{-1}$ where $k_{xy}$ is the in-plane thermal conductivity of the PGS absorber, $h$ is thickness, and $L$ is the characteristic length of heat transfer (Fig. 2E, Text S1). $R$ thus provides an useful parameter for examining the thermal response of photoabsorbers with different geometries. Varying $R$ yielded thermal response time scale values in the range 10 ms $\leq\tau\leq$ 100 ms. Pixels with higher values of $R$ yielded greater displacements, $d$, and larger values of $\tau$ (Fig. 2F). These competing effects illustrate a trade off between response amplitude and speed that we examined in further experiments.
Rapid Pixel Refresh in Cyclic Photostimulation
Achieving rapid pixel response rates is essential for meeting the perceptual requirements of tactile displays. Refresh intervals shorter than 50 ms are necessary to ensure perceptual continuity of tactile patterns across actuation cycles  (?, ?), similar to frame rate requirements for persistence of vision in visual displays. Rapid response rates also enhance the bandwidth of information conveyance and increase display responsiveness. Our display exhibits rapid actuation cycles, from 2 to 100 ms, that contrast with slower response times, 1 or more seconds, that are often obtained with thermally mediated actuation (Fig. S16). This performance is enabled by the low thermal mass of the photoabsorber ($100$ $\mu$J/K) and encapsulated air ($10$ $\mu$J/K), by the small quantity of heat energy delivered in each pulse (2.5 to 125 mJ), and by favorable heat transfer characteristics, including the surface area to volume ratio, $A/V\propto 1/L$, where $L=3$ mm is the characteristic pixel dimension.
However, achieving rapid actuation cycles is balanced by the competing need for sufficient dynamic range in force and displacement. This trade-off is parametrized by the effective thermal resistance, $R$, of the absorber. To evaluate this balance in our system, we experimentally characterized the membrane force and displacement responses across a range of optical stimuli (Fig. 3). We observed the peak output force to increase linearly with increasing optical power $P_{L}$ (Fig. 3B). At constant power, force and displacement increased monotonically with pulse length $t_{p}$ (Fig. 3B, Fig. S7). Photostimulation of higher-resistance pixels ($R=382$ K/W) yielded greater peak forces ($F=55$ mN) after $t_{p}=50$ ms, while lower resistance pixels ($R=145$ K/W) yielded smaller peak forces ($F=31$ mN) (Fig. 3B; Fig. S9).
Using these measurements, we computed the energetic efficiency, $\eta,$ of actuation as the ratio of stroke power to absorbed optical power, obtaining $\eta=0.03\%$ (Text S3, Fig. S11). This efficiency is within the range observed for many thermal actuation methods (Table S1). It reflects several losses. More than 90% is due to photoabsorber heat transfer to the walls via the bridges (Text S3). This thermal conduit speeds the pixel response time, but decreases efficiency. Adjusting $R$ allows us to tune performance, trading off efficiency and speed. Heat in the gas is converted to output work with about 14% efficiency. This loss is primarily due to entropy generation  (?), a measure of irreversibility, which is magnified at small scales. In our device, the pixel volume is $V=7.8$ $\mu$L. A theoretical scaling analysis indicates that when the characteristic dimension $L$ of a thermally actuated device is sufficiently small, the relative loss in efficiency $\eta$ scales as $\Delta\eta/\eta_{r}\propto L^{-\alpha}$, where $\eta_{r}$ is the Carnot efficiency, and $\alpha=1$ or $2$ reflects the thermodynamic design (Text S4). Similar size-efficiency effects have been experimentally observed across many thermodynamic systems  (?, ?).
Figure 3: Fig. 3. Cyclic photostimulated actuation. A) Isometric force and unloaded displacement measurements. B) Individual pixels produced peak forces of $F=$ 55 mN under photostimulation at power $P_{L}=2.5$ W and pulse duration $t_{p}=50$ ms (pulse energy 125 mJ). Force increased linearly with power $P_{L}$ and increased with pulse duration $t_{p}$. C) Successive displacement responses to pulsed excitation exhibit minimal variation for pulse gap intervals $t_{g}\gg\tau$ (Fig. S7), where $\tau$ is the thermal relaxation time; Here, $\tau\approx 23$ ms and $R=145$ K/W. At shorter intervals $t_{g},$ residual heat yields a DC offset $\delta_{o},$ reducing peak-to-peak displacement $\delta_{pp}$ relative to initial displacement $\delta_{1}$. D) Steady-state amplitude $\delta_{pp}/\delta_{1}$ (blue) increased with increasing pulse gap $t_{g}$. Blue line: $\delta_{pp}/\delta_{1}=1-\exp{(t_{g}/\tau)}$ ($r^{2}=0.95$). Normalized offset $\delta_{o}/\delta_{1}$ (red) followed an opposite trend ($r^{2}=0.81$). (Uncertainty bars = 1 SD, from $n=5$ oscillations). E) Displacements $\delta_{pp}$ for response rates up to $f=$ 500 Hz. Peak-to-peak displacement decreased as $f=1/(t_{p}+t_{g})$ increased ($t_{p}/t_{g}=1$). Inset: displacement $d(t)$ in steady-state for $f=20$ Hz and $200$ Hz. F) In sequential scanning of pixels, the maximum pixel display rate $N$ is dictated by the required displacement. At fixed $P_{L}$, $N$ decreases as displacement increases. For $P_{L}=2.5$ W, $N=$ 217 pixels/s is achieved at 50 $\mu$m, and $N=26.5$ pixels/s is achieved at 400 $\mu$m.
We next investigated the dynamic response of the display under rapid cyclic photostimulation. Rapid stimulation produces a multimodal displacement response $d(t)=\delta(t)+\delta_{o}(t)$ with slowly varying and oscillating components, $\delta(t)$ and $\delta_{o}(t)$ respectively (Fig. 3C). The slow component, $\delta(t)$, is due to the retention of heat in the cavity across cycles, as occurs when the inter-pulse gap time, $t_{g}$, is shorter than $\tau.$ This component attenuates in amplitude as $t_{g}$ increases; in steady-state, $\delta_{o}/\delta_{1}=\tau/t_{g}$ (Fig. 3D, red). The oscillating component, which is felt as a vibration, exhibits an opposite trend, described by $\delta_{pp}/\delta_{1}=1-\exp(-t_{g}/\tau)$ (Fig. 3D, blue). Across a range of values of $t_{g}$, both components fall within the perceptually salient amplitude range  (?). Thus, by adjusting the pulse gap time, $t_{g}$, we can control the relative proportion of slow and vibratory tactile feedback.
The pulse rate of cyclic photostimulation $f=1/T$ is the inverse of the pulse period, $T=t_{p}+t_{g}$. With optical power $P_{L}$ and duty cycle $\nu=t_{p}/T$ held constant, the quasi-static displacement, $\delta_{o},$ is nearly invariant with respect to variations in pulse rate $f$ (Fig. S10). This occurs because the average absorbed power $\overline{P}=\nu\varepsilon P_{L}$ is determined by the duty cycle and optical power. In contrast, because $t_{p}$ and $t_{g}$ both decrease as $f$ increases, the oscillation amplitude, $\delta_{pp},$ decreases (Fig. 3E). Nonetheless, even rapid pulsing at $f=200$ Hz yields an oscillation amplitude of $\delta_{pp}=8.4$ $\mu$m, which is large enough to be perceived via touch. For comparison, this amplitude is about 25 dB higher than the perceptual detection threshold for 200 Hz vibrations at the fingertip  (?).
While the response rate of cyclically driven pixels is constrained by the thermal response time scale, our display can also render tactile patterns by rapidly scanning the light source, at 100% duty cycle, across arrays of pixels in sequence. In sequential scanning, the maximum display rate $N$, in pixels per second, is dictated by the pulse time, $t_{p},$ needed to achieve a specified displacement amplitude, $d$, at the operating power level, $P_{L}$. The maximum pixel rate $N$ thus decreases with increasing displacement (Fig. 3F). At power $P_{L}=2.5$ W, specifying displacement to be $d=$ 50 $\mu$m results in a maximum display rate of $N=$ 217 pixels/s, while for $d=$ 400 $\mu$m the maximum rate is $N=26.5$ pixels/s.
Tactile Pattern Display and Perception
Figure 4: Fig. 4. Perception of Dynamic Tactile Pattern Display. A) Experiment 1: Perception of linear tactile motion patterns. All participants correctly reported the motion direction in 100% of trials. B) Experiment 2: Perception of rotational motion for motion speeds $v=$ 16, 32, and 64 mm/s. Mean response accuracy: 94.7%. Black dots show the overall mean, and grey dots show participant mean. C) Experiment 3: Spatial localization of actuated pixels near the finger. Mean accuracy: 78.4% (99% within 1 pixel distance). Mean localization error: 0.17 mm. D) Experiment 4: The perceived intensity of pulse train feedback from a single pixel increased linearly with optical power $P_{L}$ ($\alpha=0.022$, $\beta=-0.42$, $r^{2}=0.98$). Stimulus parameters: $N=20$ pulses, $t_{p}=t_{g}=$ 25 ms, $2.50.5<P_{L}<$ 2.5 W. E) Experiment 5: Discrimination of temporal patterns. ”Odd-One-Out” task to identify which of three sensations was different. Total accuracy: 89.6%. F) Experiment 6: Perception of multi-point tactile patterns felt at any two of four fingers. Participants reported the locations that were stimulated with 96.3% accuracy.
This tactile display employs a physical operating principle that diverges from those used in prior haptic technologies. The forces and displacements that can be produced by individual pixels (up to 55 mN and 0.97 mm) are well within the perceivable range  (?). These basic capabilities led us to investigate the potential of using multiple pixels to render a variety of perceptually distinct tactile patterns. We conducted six experiments assessing the perception of linear and rotational motion patterns, tactile localization, the scaling of perceived intensity with power, the perception of temporal patterns, and multi-finger tactile perception (Figs. 4A-F). The same ten participants completed all six experiments without training. In post-hoc surveys, no participants reported perceiving thermal sensations or seeing any optical stimuli.
Experiment 1 assessed the perception of linear motion patterns that were generated through the sequential activation of pixels along each of four cardinal directions (Fig. 4A). All participants correctly reported the motion direction in 100% of trials. In experiment 2, participants discriminated rotational motion patterns at three different motion speeds (Fig. 4B). Their response accuracy was 94.6%. These results underscore the effectiveness of the display in representing motion.
In experiment 3, participants reported which single pixel was actuated out of nine pixels arranged in a 1 cm $\times$ 1 cm grid beneath the finger pad (Fig. 4C). They responded correctly on 78.4% of trials. The mean error in localization was 0.17 mm, with 99% of all responses within one pixel of the correct location. As a point of comparison, the two point discrimination threshold, a standard measure of tactile acuity, averages 2 to 3 mm at the fingertip  (?).
In experiment 4, we assessed the relationship between perceived intensity and optical power (Fig. 4D), using the psychophysical method of magnitude estimation. Participants rated the intensity of impulse trains that were excited at different power levels (0.5 to 2.5 W). Perceived intensity increased linearly with optical power ($r^{2}=0.98$), and thus increased linearly with force (Fig. 3B). Thus, perceptually distinct levels of tactile feedback can be generated by modulating power.
In experiment 5, participants discriminated temporal patterns (Fig. 4E), using an odd-one-out paradigm. In each trial, they felt a sequence of three pulse train stimuli generated at a single pixel. Two stimuli had the same pulse rate, either slow or fast, while the third stimulus – the oddball – had a different pulse rate (respectively fast or slow). Participants correctly discriminated the temporal patterns, identifying the oddball stimulus, in 89.6% of all trials.
In experiment 6, we investigated the perception of multi-point tactile pattern perception. In each trial, participants simultaneously felt stimuli on two out of four fingers in contact with the display, and reported the pair of stimulated locations. Participants correctly identified the pair of locations with 96.3% accuracy.
Conclusion
Our tactile display technique exploits the advantages of projected light for power transmission, addressing, and actuation. Light is thermodynamically transduced into forces and displacements via energetically passive surfaces populated with optotactile pixels. The actuation mechanism, layered architecture, and simplicity of fabrication make it straightforward to produce displays with dozens to thousands of pixels, with similar investments in material cost and fabrication time. We present implementations with up to 1,511 pixels – significantly more than any previous device capable of meeting similar performance requirements. Our technique furthermore yields tactile patterns that can be not only felt, but also seen, enabling multisensory tactile displays. These displays hold promise for applications ranging from automotive interfaces to virtual reality controllers. In order for them to achieve widespread adoption, higher pixel display rates are needed. This could be achieved using optimized architectures that exploit fundamental knowledge about thermodynamic device efficiency, and by leveraging emerging high-power optical projection technologies for additive manufacturing and automotive applications  (?, ?).
Acknowledgments
We gratefully acknowledge Vijay Kumar, Yangying Zhu, Shantal Adajian, Bolin Liao, and Alban Sauret for assistance with characterization experiments. We thank Neeli Tummala for valuable comments. We acknowledge the Innovation Workshop and Microfluidics Lab in the UCSB California NanoSystems Institute, the UCSB MRSEC facility (NSF DMR 0520415), and the UCSB Nanofabrication facility, which were used in this research.
Author Contributions: M.L. and Y.V., designed research; M.L, G.R., D.G. and Y.V. performed research; M.L., E.H., and Y.V. analyzed data; M.L, and Y.V. wrote the paper.
Competing Interests: The authors declare that they have no competing interests.
Data and Materials Availability: All data needed to evaluate the conclusions in the paper are present in the paper and the Supplementary Materials.